OpenStudy (anonymous):

Given cos(2x+y)=x^2sin2y , where 0≤y≤π/2 Find dy/dx for cos(2x+y)=x^2sin2y

OpenStudy (anonymous):

implicit diff?

OpenStudy (anonymous):

oh damn i made a mistake!!!

OpenStudy (anonymous):

\[-\sin(2x+y)(2+y')=2x\sin(2y)+x^2\cos(2y)y'\] so it is even more of a pain in the butt